OpenStudy (anonymous):

Simplify.square root of 10 times square root of 8 square root of 18 4 square root of 5 square root of 80 None of the above

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

\[\sqrt{a}*\sqrt{b} = \sqrt{a*b}\]so you can get \[\sqrt{10}*\sqrt{8}=\sqrt{10*8}=\sqrt{80}\]now can you simplify that? :)

OpenStudy (anonymous):

I'm not sure how to @jigglypuff314

OpenStudy (anonymous):

Can you explain how I simplify it please?

jigglypuff314 (jigglypuff314):

for example \[\sqrt{32}=\sqrt{16*2}=\sqrt{16}*\sqrt{2}=4\sqrt{2}\]

OpenStudy (anonymous):

So is it: 4 square root of 5 ? I tried to get the right answer but I'm not sure if this is it!

OpenStudy (anonymous):

I'm not sure if it's right?

jigglypuff314 (jigglypuff314):

4 sqrt 5 is correct! :)

OpenStudy (anonymous):

Okay thanks, you get a medal!!!!

jigglypuff314 (jigglypuff314):

Glad I could help! :D